OpenStudy (anonymous):

6x+29=5 please help solve

OpenStudy (anonymous):

Will medal!

OpenStudy (anonymous):

-4

OpenStudy (uri):

6x=5-29? @owlbteen

OpenStudy (anonymous):

6x+29=5

OpenStudy (anonymous):

@uri

OpenStudy (uri):

o_o I was asking what's 5-29?

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

-24

OpenStudy (uri):

6x=-24 x=-24/6?

OpenStudy (anonymous):

-4

OpenStudy (uri):

yeah

OpenStudy (anonymous):

divide tht by 6 its -4x

OpenStudy (anonymous):

x=-4?

OpenStudy (uri):

yup.

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome

OpenStudy (uri):

yw